OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form.

OpenStudy (anonymous):

use \[y-y_1=m(x-x_1)\] with \[x_1=2,y_1=5,m=3\]